The growth in per share FCFE of CBS, Inc. is expected to be 10% per year for the next two years, followed by a growth rate of 5% per year for three years. After this five-year period, the growth in per share FCFE is expected to be 2% per year, indefinitely. The required rate of return on CBS, Inc. is 12%. Last year's per share FCFE was $2.00. What should the stock sell for today?
Answer #1

First we compute the FCFE for 6 years as below

Year FCFE
1 $2.20
2 $2.42
3 $2.54
4 $2.67
5 $2.80
6 $2.86

Horizon value = P5 = FCFE6/ (k-g)

=2.86/(12%-2%) = 28.57

Stock price today = FCF1/(1+r)^1 + FCF2/(1+r)^2 …………(FCFn+Pn)/(1+r)^n

= 2.2/1.12^1+ 2.42/1.12^2+ 2.54/1.12^3+ 2.67/1.12^4+ (2.8+28.57)/1.12^5

=$25.20